The Data of Macroeconomics-End of Chapter Problenm Abby consumes only apples. In year 1, red apples cost $1 each, green apples cost $2 each, and Abby buys 10 red apples. In year 2, red apples cost $2 each, green apples cost $1 each, and Abby buys 10 green apples. Assume that year 1 is the base year in which the consumer basket is fixed. a. Compute the CPI for apples for each year. Assume that year 1 is the base year in which the consumer basket is fixed. How does your index change from year 1 to year 2? 0 The CPI for year :100 The CPI for year 2 is: 200 CPI has doubled. - b. Compute Abbys nominal spending on apples in each year. How does it change from year 1 to year 2? Nominal Spending in year l: 10 Nominal spending in year 2: 10 Nominal spending hasstayed the same. c. Using year 1 as the base year, compute Abbys real spending on apples in each year. How does it change from year 1 to year 2? Real spending in year 1: $ 10 Real spending in year 2: $ 20 Real spending has doubled

d. Defining the implicit price deflator as nominal spending divided by real spending, compute the deflator for each year. How does the deflator change from year 1 to year 2? 0 0 Deflator in year 2: 50 The deflator has fallen by half. e. Suppose that Abby is equally happy eating red or green apples. How much has the true cost of living increased for Abby? Compare this answer to your answers to parts (a) and (d) The true cost of living has stayed the same.

Need help with the answers marked red (they are incorrect). Please don't say the cpi for year 1 is 100 (that is incorrect according to this). Please show work.
Answer #1

Ques 1

CPI for the first year will be the sum of all the apples in year 1 divided by the prices of apples in the base year (Year 1)

That is, $1 + $2 = $ 3

CPI year 1 = ( $3 ) / ( $3) x 100 = 100

CPI year 2 = ( 2 +1 ) / ( 3) x100 = 100

Hence it stayed the same

Ques d

GDP deflator = (Nominal GDP) / (Real GDP) x 100

For Year 1,

Abby buys 10 red apple @ $ 1

Nominal GDP = 10x$1 = $ 10

Real GDP(Year 1 base year) = 10 x $1 = $ 10

GDP Deflator = 100%

For Year 2,

Abby buys 10 green apple @ $ 1

Nominal GDP = 10 x $ 1 = $ 10

Real GDP( Base year 1) = 10 x 2 = $ 20

GDP Deflator = 10/20 x 100 = 50%

Fallen by 50%